how many words can be formed using letter of 'STUDENT' using each letter at most once if some or all the letters may be omitted
Answer by AnlytcPhil(1806)

I will assume the two T's are indistinguishable and 
that there can be two T's in the same word.  

The number of words that have all different letters is

6P1 + 6P2 + 6P3 + 6P4 + 6P5 + 6P6 = 

6 + 30 + 120 + 360 + 720 + 720 = 1956

In addition to that number, we must calculate:

A. The number of 2 letter words with 2 T's
There's only 1, which is TT.

B. The number of 3 letter distinguishable words with 2 T's:
There are 3C2 = 3 positions to place the 2 T's. 
There are 5P1 = 5 ways to place the 1 non-T.
That's 3*5 = 15 ways. 

C. The number of 4 letter distinguishable words with 2 T's:
There are 4C2 = 6 positions to place the 2 T's. 
There are 5P2 = 20 ways to place the 2 non-T's.
That's 6*20 = 120 ways.

D. The number of 5 letter distinguishable words with 2 T's:
There are 5C2 = 10 positions to place the 2 T's. 
There are 5P3 = 60 ways to place the 3 non-T's.
That's 10*60 = 600 ways.

E. The number of 6 letter distinguishable words with 2 T's:
There are 6C2 = 15 positions to place the 2 T's. 
There are 5P4 = 120 ways to place the 4 non-T's.
That's 15*120 = 1800 ways.

F. The number of 7 letter distinguishable words with 2 T's:
There are 7C2 = 21 positions to place the 2 T's. 
There are 5P5 = 120 ways to place the 5 non-T's.
That's 21*120 = 2520 ways.

[As a check on that last number of ways, we can use the usual 
formula for the number of distinguishable words that can be 
formed from a 7-letter word with 2 indistinguishable letters, 
7!/2! = 5040/2 = 2520]

The total number of words with two T's is
 
1 + 15 + 120 + 600 + 1800 + 2520 = 5056

Grand total = 1956 + 5056 = 7012 

Edwin